OpenStudy (anonymous):

what is the square root of 8x3

OpenStudy (luigi0210):

This? \[\LARGE \sqrt{8*3}\]

OpenStudy (luigi0210):

If that's it, just take it apart.. 8=4*2 \[\LARGE \sqrt{4*2*3}\] Since 4 is a perfect square, you can take it out and are left with 2*3. That equals 6, and we can't do anything with that so we just leave it as is: \[\LARGE 2\sqrt{6}\]